# Matbench Discovery A framework to evaluate machine learning crystal stability predictions

### v2

@@ 1.1.2 WBM Test Set
The WBM data set [29] consists of 257 k structures generated via chemical similarity-based elemental substitution of MP source structures followed by DFT relaxation and calculating each crystal’s convex hull distance. Which element replaces an existing one in a given source structure was determined by random sampling according to the weights in a chemical similarity matrix data-mined from the ICSD [33].
@@
The WBM authors performed 5 iterations of this substitution process (we refer to these as batches). After each step, the newly generated structures found to be thermodynamically stable after DFT relaxation flow back into the source pool to partake in the next round of substitution. This split of the data into batches of increasing substitution count is a unique and compelling feature of the test set as it allows out-of-distribution (OOD) testing by seeing if model performance degrades with substitution count. A higher number of elemental substitutions on average carries the structure further away from the region of material space covered by the MP training set (see figs. 8 and 10 for details). Whilst this batch information makes the WBM data set an exceptionally useful data source for examining the extrapolation performance of ML models, our evaluation primarily looks at metrics that consider all batches as a single test set.
@@
Throughout this work, we define stability as being on or below the convex hull of the MP training set (). 42k out of 257k materials in WBM satisfy this criterion. Of these, 33k are unique prototypes, meaning they have no matching structure prototype in MP nor another higher-energy duplicate prototype in WBM. Our code treats the stability threshold as a dynamic parameter, allowing for future model comparisons at different thresholds. This may reveal systematic differences in stability calibration, as hinted by the different TPR/TNR tradeoffs observed for CHGNet and M3GNet in table 3. A looser stability threshold (e.g. ) might benefit a model like M3GNet which is more conservative in its stability predictions, whereas a stricter threshold may benefit a model like CHGNet that is more optimistic in its stability predictions. For initial analysis in this direction, see fig. 3 in the SI.
@@
As WBM explores regions of materials space not well sampled by MP, many of the discovered materials that lie below MP’s convex hull are not stable relative to each other. Of the 42k that lie below the MP convex hull less than half, or around 20k, remain on the joint MP+WBM convex hull. This observation suggests that many WBM structures are repeated samples in the same new chemical spaces. It also highlights a critical aspect of this benchmark in that we purposely operate on an incomplete convex hull. Only current knowledge of competing points on the PES is accessible to a real discovery campaign and our metrics are designed to reflect this.
@@ 1.2 Models
To test a wide variety of methodologies proposed for learning the potential energy landscape, our initial benchmark release includes 10 models.
@@
MACE [20] (UIP-GNN) - MACE builds upon the recent advances [34, 35] in equivariant neural network architectures by proposing an approach to computing high-body-order features efficiently via Atomic Cluster Expansion [36]. Unlike the other UIP models considered MACE was primarily developed for molecular dynamics of single material systems and not the universal use case studied here. It is the only -equivariant model we tested.
@@
CHGNet [19] (UIP-GNN) - CHGNet is a UIP for charge-informed atomistic modeling. Its distinguishing feature is that it was trained to predict magnetic moments on top of energies, forces and stresses in the MPTrj dataset consisting of relaxation trajectories for 1.5 million MP structures (see section A.10 for detailed analysis of this dataset). By modeling magnetic moments, CHGNet learns to accurately represent the orbital occupancy of electrons which allows it to predict both atomic and electronic degrees of freedom.
@@
M3GNet [18] (UIP-GNN) - M3GNet is a GNN-based universal interatomic potential (UIP) (as in covering the full periodic table) for materials trained on up to 3-body interactions in the initial, middle and final frame of MP DFT relaxations. The model takes the unrelaxed input and emulates structure relaxation before predicting energy for the pseudo-relaxed structure.
@@
ALIGNN [37] (GNN) - The Atomistic Line Graph Neural Network (ALIGNN) is a message passing GNN architecture that takes as input both the interatomic bond graph and a line graph corresponding to 3-body bond angles. The ALIGNN architecture involves a global pooling operation which means that it is ill-suited to force-field applications. To address this the ALIGNN-FF model was later introduced without global pooling [38].
@@
MEGNet [39] (GNN) - MatErials Graph Network is another GNN similar to CGCNN for material properties of relaxed structures that also updates the edge and global features (like pressure, temperature, entropy) in its message passing operation. This work showed that learned element embeddings encode periodic chemical trends and can be transfer-learned from large data sets (formation energies) to predictions on small data properties (band gaps, elastic moduli).
@@
CGCNN [28] (GNN) - The Crystal Graph Convolutional Neural Network (CGCNN) was the first neural network model to directly learn 8 different DFT-computed material properties from a graph representing the atoms and bonds in a periodic crystal. CGCNN was among the first to show that just like in other areas of ML, given large enough training sets, neural networks can learn embeddings that outperform human-engineered structure features directly from the data.
@@
CGCNN+P [40] (GNN) - This work proposes simple, physically motivated structure perturbations to augment stock CGCNN’s training data of relaxed structures with structures resembling unrelaxed ones but mapped to the same DFT final energy. Here we chose , meaning the training set is augmented with 5 random perturbations of each relaxed MP structure mapped to the same target energy.
In contrast to all other structure-based GNNs considered in this benchmark, CGCNN+P is not attempting to learn the Born-Oppenheimer potential energy surface. The model is instead taught the PES as a step-function that maps each valley to its local minimum. The idea is that during testing on unrelaxed structures, the model will predict the energy of the nearest basin in the PES. The authors confirm this by demonstrating a lowering of the energy error on unrelaxed structures. -
@@
Wrenformer (Transformer) - For this benchmark, we introduce Wrenformer which is a variation on the coordinate-free Wren model [10] constructed using standard QKV-self-attention blocks [41] in place of message-passing layers. This architectural adaptation reduces the memory usage allowing the architecture to scale to structures with greater than 16 Wyckoff positions.
Like its predecessor, Wrenformer is a fast coordinate-free model aimed at accelerating screening campaigns where even the unrelaxed structure is a priori unknown. The key idea is that by training on the coordinate anonymized Wyckoff positions (symmetry-related positions in the crystal structure), the model learns to distinguish polymorphs while maintaining discrete and computationally enumerable inputs. The central methodological benefit of an enumerable input is that it allows users to predict the energy of all possible combinations of spacegroup and Wyckoff positions for a given composition and maximum unit cell size. The lowest-ranked prototypes can then be fed into downstream analysis or modeling. -
@@
BOWSR [17] (BO-GNN) - BOWSR combines a symmetry-constrained Bayesian optimizer (BO) with a surrogate energy model to perform an iterative exploration-exploitation-based search of the potential energy landscape. Here we use MEGNet [39] for the energy model as proposed in the original work. The high sample count needed to explore the PES with BO makes this by far the most expensive model tested.
@@
Voronoi RF [42] (Fingerprint) - A random forest trained to map a combination of composition-based Magpie features [43] and structure-based relaxation-robust Voronoi tessellation features (effective coordination numbers, structural heterogeneity, local environment properties, …) to DFT formation energies.
This fingerprint-based model predates most deep learning for materials but significantly improved over earlier fingerprint-based methods such as the Coulomb matrix [44] and partial radial distribution function features [45]. It serves as a baseline model to see how much value the learned featurization of deep learning models can extract from the increasingly large corpus of available training data.
@@ 2 Results
Table 1 shows performance metrics for all models included in the initial release of Matbench Discovery. MACE takes the top spot and emerges as the current SOTA for ML-guided materials discovery. It outperforms the 9 other models on all 6 classification metrics but is slightly overtaken by CHGNet in 2 of 3 regression metrics, and RMSE, which are sensitive to outliers. The discovery acceleration factor (DAF) measures how many more stable structures a model found compared to the dummy discovery rate achieved by randomly selecting test set crystals, formally the DAF is the ratio of the precision to the prevalence. The maximum possible DAF is the inverse of the prevalence or dummy discovery rate which on our dataset is . Thus the current SOTA of 3.85 achieved by MACE leaves room for improvement. However, evaluating each model on the much smaller subset of its 10k most stable predictions as shown in table 3, CHGNet takes the lead in all 9 metrics, in particular achieving an impressive DAF of 5.29 which approaches optimal performance.
@@
We find a large performance gap between models that make one-shot predictions directly from unrelaxed inputs (MEGNet, Wrenformer, CGCNN, CGCNN+P, ALIGNN, Voronoi RF) compared to UIPs that learn from force and stress labels to emulate DFT relaxation (CHGNet, M3GNet, MACE). While the F1 scores and DAFs of force-free models are surprisingly less affected, their regression metrics such as and RMSE are significantly worse. Of the force-free models, only ALIGNN, BOWSR and CGCNN+P achieve positive predictive power . Negative means model predictions explain the observed variation in the data less than simply predicting the test set mean. In other words, these models are not predictive in a global sense (across the full dataset range). However, even models with negative can be locally good in the positive and negative tails of the test set distribution. They suffer most in the mode of the distribution near the stability threshold of above the hull. This reveals an important shortcoming of as a metric for classification tasks like stability prediction.
The results for M3GNet depart from the trend that F1 is broadly rank-correlated with the other classification metrics. M3GNet achieves a high true positive rate () but a notably lower true negative rate () than CHGNet and ALIGNN with the closest F1 scores. Figure 2 provides a visual explanation for this observation. M3GNet has a lower rolling mean of the absolute errors (rolling MAE) as a function of hull distance for materials above the convex hull (see right half of plot) but incurs comparably large errors for materials below the hull (left half of plot). Since , lower error for energies above the hull increases both TN and decreases FP, resulting in the high TPR values observed.
@@
The reason CGCNN+P achieves better regression metrics than CGCNN but is still worse as a classifier becomes apparent from fig. 7 by noting that the CGCNN+P histogram is more sharply peaked at the 0 hull distance stability threshold. This causes even small errors in the predicted convex hull distance to be large enough to invert a classification. Again, this is evidence to choose carefully which metrics to optimize. Regression metrics are far more prevalent when evaluating energy predictions. However, our benchmark treats energy predictions as merely means to an end to classify compound stability. Improvements in regression accuracy are of limited use to materials discovery in their own right unless they also improve classification accuracy. Our results demonstrate that this is not a given.
@@
Figure 1 has models rank materials by model-predicted hull distance from most to least stable; materials furthest below the known hull at the top, materials right on the hull at the bottom. For each model, we iterate through that list and calculate at each step the precision and recall of correctly identified stable materials. This simulates exactly how these models would be used in a prospective materials discovery campaign and reveals how a model’s performance changes as a function of the discovery campaign length. As a practitioner, you have a certain amount of resources available to validate model predictions. These curves allow you to read off the best model given these conditions and based on the optimal trade-off between fewer false positives (precision) or fewer negatives (recall) for the discovery task at hand. In this case, it so happens that CHGNet achieves the highest precision and recall at any number of screened materials.
@@
A line terminates when a model believes there are no more materials in the WBM test set below the MP convex hull. The dashed vertical line shows the actual number of stable structures in our test set. All models are biased towards stability to some degree as they all overestimate this number, most of all BOWSR by 133%, least of all MEGNet by 30%. This is only a problem in practice for exhaustive discovery campaigns that validate all stable predictions from a model. More frequently, model predictions will be ranked most-to-least stable and validation stops after some pre-determined compute budget is spent, say, 10k DFT relaxations. In that case, the concentration of false positive predictions that naturally accumulates near the less stable end of the candidate list can be avoided with no harm to the campaign’s overall discovery rate (see table 3 a similar metrics table considering only the 10k materials predicted by each model to be furthest below the known convex hull).
The diagonal Optimal Recall line would be achieved if a model never made a false negative prediction and stopped predicting stable crystals exactly when the true number of stable materials is reached. Zooming in on the top-left corner of the precision plot, we observe that CHGNet is the only model without a sudden drop in precision right at the start of the discovery campaign. This means CHGNet is the only model whose first few hundred most stable materials are largely actually stable. MACE in particular suffers from several initial failure cases. MACE underpredicts the energy of these unstable materials by several eV/atom (see fig. 4), resulting in MACE’s initial precision starting at 0 before abruptly climbing to 0.8. We experienced this behavior with multiple independently trained MACE variants; even more so with a checkpoint we received directly from the MACE authors which was trained on the M3GNet training set. The results shown here are from a superior MACE we trained ourselves on the much larger MPtrj dataset. Even M3GNet, the other universal potential and 2nd best model, exhibits the early-on precision drop. As a result, CHGNet has a strong lead over M3GNet until reaching 3k screened materials. From there, CHGNet and M3GNet slowly converge until they almost tie at a precision of 0.52 after 56k screened materials. At that point, CHGNet’s list of stable predictions is exhausted while M3GNet continues, dropping in precision to 0.45 at 76 k, attributable to many false positives near the end of the list of stable predictions.
@@
All force-free models exhibit a much worse case of early-on precision drop, falling to 0.6 or less in the first 5k screened materials. Many of these models (all except BOWSR, Wrenformer and Voronoi RF) display an interesting hook shape in their cumulative precision, recovering again slightly in the middle of the simulated campaign between 5k and up to 30k before dropping again until the end.
@@
Figure 2 provides a visual representation of the reliability of different models based on the rolling mean absolute error (MAE) of model-predicted hull distances as a function of DFT distance to the Materials Project (MP) convex hull. The red-shaded area, which we coin the ‘triangle of peril’, emphasizes the zone where the average model error surpasses the distance to the stability threshold at 0 eV. As long as the rolling MAE remains within this triangle, the model is most susceptible to misclassifying structures. Because the average error is larger than the distance to the classification threshold at 0, it is large enough to flip a correct classification into an incorrect one (if the error happens to point toward the stability threshold).
The sooner a model leaves the triangle on the left side, the less likely it is to incorrectly predict stable structures as unstable, thereby reducing false negatives. On the right side, early exits result in a lower likelihood of predicting unstable structures as stable, leading to fewer false positives.
On the left side, CHGNet exits the triangle first followed closely by ALIGNN. We can expect fewer false negatives from CHGNet than from other models. On the right side, CHGNet and M3GNet exit simultaneously, at a remarkably low error, followed shortly after by MACE at . All 3 UIPs therefore produce fewer false positives than other models. For highly unstable structures, first CHGNet and later M3GNet are both overtaken by Wrenformer for highly unstable structures towards the right end of the plot.
Overall, this visualization underscores the importance of considering a model’s average error relative to the distance to the stability threshold when evaluating its performance as a material stability predictor. Low overall hull distance error can be a misleading metric if that error is not particularly low at the decision boundary.
The imbalance between the left and right half of the plot shows that models are more prone to false negative predictions, even for very stable materials far below the known hull, than predicting very unstable materials as stable.
As the convex hull becomes more thoroughly sampled by future discovery, the fraction of unknown stable structures decreases, naturally leading to less enriched future test sets. This has several implications, including allowing for higher maximum DAFs but also skewing the hull distance distribution towards positive values (i.e. the right half of fig. 2) as there are fewer undersampled chemical spaces. Consequently, to have accurate stability classifications, model accuracy will need to improve concomitantly. For fig. 2, this means models need to be much more accurate to exit the shaded triangle in the left half of the plot.

### v3

@@ 2.1.2 WBM Test Set
The WBM dataset [wang_predicting_2021] consists of 257,487 structures generated via chemical similarity-based elemental substitution of MP source structures followed by DFT relaxation and calculating each crystal’s convex hull distance. The element substitutions applied to a given source structure were determined by random sampling according to the weights in a chemical similarity matrix data-mined from the ICSD [glawe_optimal_2016].
@@
The WBM authors performed 5 iterations of this substitution process (we refer to these steps as batches). After each step, the newly generated structures found to be thermodynamically stable after DFT relaxation flow back into the source pool to partake in the next round of substitution. This split of the data into batches of increasing substitution count is a unique and compelling feature of the test set as it allows out-of-distribution (OOD) testing by examining whether model performance degrades for later batches. A higher number of elemental substitutions on average carries the structure further away from the region of material space covered by the MP training set (see fig. S6 for details). Whilst this batch information makes the WBM dataset an exceptionally useful data source for examining the extrapolation performance of ML models, we look primarily at metrics that consider all batches as a single test set.
In order to control for the potential adverse effects of leakage between the MP training set and the WBM test set, we cleaned the WBM test set based on protostructure matching. We refer to the combination of a materials prototype and the elemental assignment of its wyckoff positions as a protostructure following [parackal2024identifying]. First we removed 524 pathological structures in WBM based on formation energies being larger than 5 eV/atom or smaller than -5 eV/atom. We then removed from the WBM test set all examples where the final protostructure of a WBM material matched the final protostructure of an MP material. In total 11,175 materials were cleaned using this filter. We further removed all duplicated protostructures within WBM, keeping the lowest energy structure in each instance, leaving 215,488 structures in the unique prototype test set.
@@
Throughout this work, we define stability as being on or below the convex hull of the MP training set (). 32,942 out of 215,488 materials in WBM unique prototype test set satisfy this criterion. Of these, 33k are unique prototypes, meaning they have no matching structure prototype in MP nor another higher-energy duplicate prototype in WBM. Our code treats the stability threshold as a dynamic parameter, allowing for future model comparisons at different thresholds. For initial analysis in this direction, see fig. S1 in the SI.
@@
As WBM explores regions of materials space not well sampled by MP, many of the discovered materials that lie below MP’s convex hull are not stable relative to each other. Of the 33k that lie below the MP convex hull less than half, or around 20k, remain on the joint MP+WBM convex hull. This observation suggests that many WBM structures are repeated samples in the same new chemical spaces. It also highlights a critical aspect of this benchmark in that we knowingly operate on an incomplete convex hull. Only current knowledge of competing points on the PES is accessible to a real discovery campaign and our metrics are designed to reflect this.
2.1.3 Limitations of this Framework
Whilst the framework proposed here directly mimics a common computational materials discovery workflow, it is worth highlighting that there still exist significant limitations to these traditional computational workflows that can prevent the material candidates suggested by such a workflow from being able to be synthesized in practice. For example, high-throughput DFT calculations often use small unit cells which can lead to artificial orderings of atoms. The corresponding real material may be disordered due to entropic effects that cannot be captured in the zero-kelvin thermodynamic convex hull approximated by DFT [cheetham2024artificial].
Another issue is that, when considering small unit cells, the DFT relaxations may get trapped at dynamically unstable saddle points in the true PES. This failure can be detected by calculating the phonon spectra for materials predicted to be stable. However, the cost of doing so with DFT is often deemed prohibitive for high-throughput screening. The lack of information about the dynamic stability of nominally stable materials in the WBM test set prevents this work from considering this important criteria as an additional screening filter. However, recent progress in the development of UIPs suggests that ML approaches will soon provide sufficiently cheap approximations of these terms for high-throughput searches [batatia2023foundation, deng2024overcoming]. As the task presented here begins to saturate, we believe that future discovery benchmarks should extend upon the framework proposed here to also incorporate criteria based on dynamic stability.
[…]
We design the benchmark considering a positive label for classification as being on or below the convex hull of the MP training set. An alternative formulation would be to say that materials in WBM that are below the MP convex hull but do not sit on the combined MP+WBM convex hull are negatives. The issue with such a design is that it involves unstable evaluation metrics. If we consider the performance against the final combined convex hull rather than the initial MP convex hull, then each additional sample considered can retroactively change whether or not a previous candidate would be labeled as a success as it may no-longer sit on the hull. Since constructing the convex hull is computationally expensive, this path dependence makes it impractical to evaluate cumulative precision metrics (see Figure 2). The chosen setup does increase the number of positive labels and could consequently be interpreted as overestimating the performance. This overestimation decreases as the convex hull becomes better sampled. Future benchmarks building on this work could make use of combination of MP+WBM to control this artifact. An alternative framework could report metrics for each WBM batch in turn and retrain between batches, this approach was undesirable here as it increases the cost of submission five-fold and introduces many complexities, for example, should each model only retrain on candidates it believed to be positive, that would make fair comparison harder.
@@ 2.2 Models
To test a wide variety of methodologies proposed for learning the potential energy landscape, our initial benchmark release includes 13 models. Next to each model’s name we give the training targets that were used: E - Energy, F - Forces, S - Stresses and M - Magnetic moments. The subscripts G and D refer to whether gradient-based or direct prediction methods were used to obtain force and stress predictions.
@@
EquiformerV2 + DeNS [liao2024equiformerV2, liao2024generalizing] (EFSD) - EquiformerV2 builds on the first Equiformer model [liao2023equiformer] by replacing the convolutions with eSCN convolutions [passaro2023escn] as well as a range of additional tweaks to make better use of the ability to scale to higher using eSCN convolutions. EquiformerV2 uses direct force prediction rather than taking the forces as the derivative of the energy predictions for computational efficiency. Here we take the pre-trained "eqV2 S DeNS" [barroso2024open] trained on the MPTrj dataset. This model in addition to supervised training on energies, forces, and stresses makes use of a auxilary training task based on de-noising non-equilibrium structures [liao2024generalizing]. We refer to this model as "EquiformerV2 + DeNS" in the text and "eqV2 S DeNS" in plots.
@@
Orb [neumann2024orb] (EFSD) - Orb is a lightweight model architecture developed to scale well for the simulation of large systems such as metal organic frameworks. Rather than constructing an architecture that is equivariant by default, Orb instead makes use of data augmentation during training to achieve approximate equivariance. This simplifies the architecture allowing for faster inference. We report results for the "orb-mptrj-only-v2" model which was pre-trained using a diffusion-based task on MPTrj before supervised training on the energies, forces and stresses in MPTrj. For simplicity we refer to this model as "ORB MPTrj".
@@
SevenNet [park2024scalable] (EFSG) - SevenNet emerged from an effort to improve the performance of message passing neural networks [gilmer2017neural] when used to conduct large scale simulations involving that benefit from parallelism via spatial decomposition. Here we use the pre-trained "SevenNet-0_11July2024" trained on the MPTrj dataset. The SevenNet-0 model is an equivariant architecture based on a NequIP [batzner_equivariant_2022] architecture that mostly adopts the GNoME [merchant2023scaling] hyper-parameters. SevenNet-0 differs from NequIP and GNoME by replacing the tensor product in the self-connection layer with a linear layer applied directly to the node features, this reduces the number of parameters from 16.24 million in GNoME to 0.84 million for SevenNet-0. For simplicity we refer to this model as "SevenNet".
@@
MACE [batatia_mace_2023] (EFSG) - MACE builds upon the recent advances [thomas_tensor_2018, batzner_equivariant_2022] in equivariant neural network architectures by proposing an approach to computing high-body-order features efficiently via Atomic Cluster Expansion [drautz_atomic_2019]. Unlike the other UIP models considered MACE was primarily developed for molecular dynamics of single material systems and not the universal use case studied here. The authors trained MACE on the MPTrj dataset, these models have been shared under the name "MACE-MP-0" [batatia2023foundation] and we report results for the "2023-12-03" version commonly called "MACE-MP-0 (medium)". For simplicity we refer to this model as "MACE".
@@
CHGNet [deng_chgnet_2023] (EFSGM) - CHGNet is a UIP for charge-informed atomistic modeling. Its distinguishing feature is that it was trained to predict magnetic moments on top of energies, forces and stresses in the MPtrj dataset (which was prepared for the purposes of training CHGNet). By modeling magnetic moments, CHGNet learns to accurately represent the orbital occupancy of electrons which allows it to predict both atomic and electronic degrees of freedom. We make use of the pre-trained "v0.3.0" CHGNet model from [deng_chgnet_2023].
@@
M3GNet [chen_universal_2022] (EFSG) - M3GNet is a GNN-based UIP for materials trained on up to 3-body interactions in the initial, middle and final frame of MP DFT relaxations. The model takes the unrelaxed input and emulates structure relaxation before predicting energy for the pseudo-relaxed structure. We make use of the pre-trained "v2022.9.20" M3GNet model from [chen_universal_2022] trained on the compliant MPF.2021.2.8 dataset.
@@
ALIGNN [choudhary_atomistic_2021] (E) - The Atomistic Line Graph Neural Network (ALIGNN) is a message passing GNN architecture that takes as input both the interatomic bond graph and a line graph corresponding to 3-body bond angles. The ALIGNN architecture involves a global pooling operation which means that it is ill-suited to force-field applications. To address this the ALIGNN-FF model was later introduced without global pooling [choudhary_unified_2023]. We trained ALIGNN on the MP-crystals-2022.10.28 dataset for this benchmark.
@@
MEGNet [chen_graph_2019] (E) - MatErials Graph Network is another GNN-based architecture that also updates a set of edge and global features (like pressure and temperature) in its message passing operation. This work showed that learned element embeddings encode periodic chemical trends and can be transfer-learned from large datasets (formation energies) to predictions on small data properties (band gaps, elastic moduli). We make use of the pre-trained "Eform_MP_2019" MEGNet model trained on the compliant MP-crystals-2019.4.1 dataset.
@@
CGCNN [xie_crystal_2018] (E) - The Crystal Graph Convolutional Neural Network (CGCNN) was the first neural network model to directly learn 8 different DFT-computed material properties from a graph representing the atoms and bonds in a periodic crystal. CGCNN was among the first to show that just like in other areas of ML, given large enough training sets, neural networks can learn embeddings that outperform human-engineered structure features directly from the data. We trained an ensemble of 10 CGCNN models on the MP-crystals-2022.10.28 dataset for this benchmark.
@@
CGCNN+P [gibson_data-augmentation_2022] (E) - This work proposes simple, physically motivated structure perturbations to augment stock CGCNN’s training data of relaxed structures with structures resembling unrelaxed ones but mapped to the same DFT final energy. Here we chose , meaning the training set is augmented with 5 random perturbations of each relaxed MP structure mapped to the same target energy. In contrast to all other structure-based GNNs considered in this benchmark, CGCNN+P is not attempting to learn the Born-Oppenheimer potential energy surface. The model is instead taught the PES as a step-function that maps each valley to its local minimum. The idea is that during testing on unrelaxed structures, the model will predict the energy of the nearest basin in the PES. The authors confirm this by demonstrating a lowering of the energy error on unrelaxed structures. We trained an ensemble of 10 CGCNN+P models on the MP-crystals-2022.10.28 dataset for this benchmark.
11.
Wrenformer (E) - For this benchmark, we introduce Wrenformer which is a variation on the coordinate-free Wren model [goodall_rapid_2022] constructed using standard QKV-self-attention blocks [vaswani_attention_2017] in place of message-passing layers. This architectural adaptation reduces the memory usage allowing the architecture to scale to structures with greater than 16 Wyckoff positions. Like its predecessor, Wrenformer is a fast coordinate-free model aimed at accelerating screening campaigns where even the unrelaxed structure is a priori unknown [parackal2024identifying]. The key idea is that by training on the coordinate anonymized Wyckoff positions (symmetry-related positions in the crystal structure), the model learns to distinguish polymorphs while maintaining discrete and computationally enumerable inputs. The central methodological benefit of an enumerable input is that it allows users to predict the energy of all possible combinations of spacegroup and Wyckoff positions for a given composition and maximum unit cell size. The lowest-ranked protostructures can then be fed into downstream analysis or modeling. We trained an ensemble of 10 Wrenformer models on the MP-crystals-2022.10.28 dataset for this benchmark.
12.
BOWSR [zuo_accelerating_2021] (E) - BOWSR combines a symmetry-constrained Bayesian optimizer (BO) with a surrogate energy model to perform an iterative exploration-exploitation-based search of the potential energy landscape. Here we use the pre-trained "Eform_MP_2019" MEGNet model [chen_graph_2019] for the energy model as proposed in the original work. The high sample count needed to explore the PES with BO makes this by far the most expensive model tested.
13.
Voronoi RF [ward_including_2017] (E) - A random forest trained to map a combination of composition-based Magpie features [ward_general-purpose_2016] and structure-based relaxation-robust Voronoi tessellation features (effective coordination numbers, structural heterogeneity, local environment properties, …) to DFT formation energies. This fingerprint-based model predates most deep learning for materials but significantly improved over earlier fingerprint-based methods such as the Coulomb matrix [rupp_fast_2012] and partial radial distribution function features [schutt_how_2014]. It serves as a baseline model to see how much value the learned featurization of deep learning models can extract from the increasingly large corpus of available training data. We trained Voronoi RF on the MP-crystals-2022.10.28 dataset for this benchmark.
@@ 3 Results
Table 1 shows performance metrics for all models included in the initial release of Matbench Discovery reported on the unique protostructure subset. EquiformerV2 + DeNS takes the top spot and emerges as the current SOTA for ML-guided materials discovery. It outperforms the other models on all 9 reported metrics. When computing metrics in the presence of missing values or obviously pathological predictions (error of 5 eV/atom or greater) we assign the dummy regression values and a negative classification prediction to these points. The discovery acceleration factor (DAF) measures how many more stable structures a model found compared to the dummy discovery rate achieved by randomly selecting test set crystals. Formally the DAF is the ratio of the precision to the prevalence. The maximum possible DAF is the inverse of the prevalence which on our dataset is . Thus the current SOTA of 5.04 achieved by EquiformerV2 + DeNS leaves room for improvement. However, evaluating each model on the subset of 10k materials each model ranks as being most stable (see table S2), we see an impressive DAF of 6.33 for EquiformerV2 + DeNS which is approaching optimal performance for this task.
@@
We find a large performance gap between models that make energy-only predictions directly from unrelaxed inputs (MEGNet, Wrenformer, CGCNN, CGCNN+P, ALIGNN, Voronoi RF) compared to UIPs that learn from force and stress labels and are used to emulate DFT relaxation before predicting a final energy. While the F1 scores and DAFs of energy-only models are surprisingly performant, their regression metrics such as and RMSE are significantly worse. Of the energy-only models, only ALIGNN, BOWSR and CGCNN+P achieve positive coefficients of determination, . Negative means model predictions explain the observed variation in the data less than simply predicting the test set mean. In other words, these models are not predictive in a global sense (across the full dataset range). However, even models with negative can be locally good in the positive and negative tails of the test set distribution. They suffer most in the mode of the distribution near the stability threshold of above the hull. This reveals an important shortcoming of as a metric for classification tasks like stability prediction.
@@
The reason CGCNN+P achieves better regression metrics than CGCNN but is still worse as a classifier becomes apparent from fig. S5 by noting that the CGCNN+P histogram is more sharply peaked at the 0 hull distance stability threshold. This causes even small errors in the predicted convex hull distance to be large enough to invert a classification. Again, this is evidence to choose carefully which metrics to optimize. Regression metrics are far more prevalent when evaluating energy predictions. However, our benchmark treats energy predictions as merely means to an end to classify compound stability. Improvements in regression accuracy are of limited use to materials discovery in their own right unless they also improve classification accuracy. Our results demonstrate that this is not a given.
@@
Figure 2 has models rank materials by model-predicted hull distance from most to least stable; materials furthest below the known hull at the top, materials right on the hull at the bottom. For each model, we iterate through that list and calculate at each step the precision and recall of correctly identified stable materials. This simulates exactly how these models would be used in a prospective materials discovery campaign and reveals how a model’s performance changes as a function of the discovery campaign length. As a practitioner, you typically have a certain amount of resources available to validate model predictions. These curves allow you to read off the best model given these constraints. For instance, plotting the results in this manner shows that CHGNet initially achieves higher precision, than models such as EquiformerV2+DeNS, ORB MPTrj, SevenNet, and MACE that report higher precision across the whole test set.
@@
In Figure 2 each line terminates when the model believes there are no more materials in the WBM test set below the MP convex hull. The dashed vertical line shows the actual number of stable structures in our test set. All models are biased towards stability to some degree as they all overestimate this number, most of all BOWSR by 133%. This is only a problem in practice for exhaustive discovery campaigns that validate all stable predictions from a model. More frequently, model predictions will be ranked most-to-least stable and validation stops after some pre-determined compute budget is spent, say, 10k DFT relaxations. In that case, the concentration of false positive predictions that naturally accumulates near the less stable end of the candidate list can be avoided with no harm to the campaign’s overall discovery rate (see table S2 where even the DAF of the worst performing model benchmarked, Voronoi RF, jumps from 1.58 to 2.49).
The diagonal ‘Optimal Recall’ line on the recall plot in Figure 2 would be achieved if a model never made a false negative prediction and stopped predicting stable crystals exactly when the true number of stable materials is reached. Examining the UIP models, we find that they all achieve similar recall values, ranging from approximately 0.75 to 0.86.. This is substantially smaller than the variation we see in the precision for the same models 0.44-0.77. Inspecting the overlap, we find that the intersection of the models’ correct agreements accounts for a precision of only 0.57 within the 0.75-0.86 range, with just 0.04 of the examples where all models are wrong simultaneously. These results indicate that the models are making meaningfully different predictions.
@@
Examining the precision plot in Figure 2, we observe that the energy-only models exhibit a much more pronouced drop in their precision early-on, falling to 0.6 or less in the first 5k screened materials. Many of these models (all except BOWSR, Wrenformer and Voronoi RF) display an interesting hook shape in their cumulative precision, recovering again slightly in the middle of the simulated campaign between 5k and up to 30k before dropping again until the end.
@@
Figure 3 provides a visual representation of the reliability of different models as a function of a material’s DFT distance to the MP convex hull. The lines show the rolling mean absolute error (RMAE) of model-predicted hull distances vs DFT. The red-shaded area, which we coin the ‘triangle of peril’, emphasizes the zone where the average model error surpasses the distance to the stability threshold at 0 eV. As long as the rolling MAE remains within this triangle, the model is highly susceptible to misclassifying structures. The average error in this region is larger than the distance to the classification threshold at 0, consequently in cases where the error points towards the stability threshold it would be large enough to flip a correct classification into an incorrect one. The sooner a model exits the triangle on the left side (negative DFT hull distance), the less likely it is to incorrectly predict stable structures as unstable, thereby reducing false negatives. Exciting early on the right side (positive DFT hull distance) results in a lower likelihood of predicting unstable structures as stable, leading to fewer false positives.
All models tend to have lower rolling error towards the plot’s left edge compared to the right edge. This imbalance shows that models are more prone to false positive than false negative predictions. In other words, all models are less likely to predict a material at DFT hull distance as unstable than a material at DFT hull distance as stable. From a practitioner’s standpoint, this is undesirable as attempting to validate an unstable material usually has a much higher opportunity cost than missing a stable one. We hypothesize this error imbalance is due to the MP training set having an uncharacteristically high share of stable materials, causing statical models trained on it to be biased towards making low-energy predictions even for high-energy atomic configurations. Training on datasets with more high-energy structures, such as Alexandria [schmidt2024improving] and OMat24 [barroso2024open], would be expected to improve performance by balancing out this source of bias.